ln 0.233 =(-1.455) but how?
1
Expert's answer
2012-09-14T10:24:52-0400
ln can be negative, because ln - it's exponentiation. { Ln 0.233=log e (0.233) }
- number needed to bring e to 0.233. But 0.233<1, so index will be negative.
(For example, log 8 (1/2) = -1/3, because 8^(-1/3)=1/2 )